# SEACrowd: A Multilingual Multimodal Data Hub and Benchmark Suite for Southeast Asian Languages

### v1

@@ 5.2 Conclusion & Future Work
Southeast Asia is home to highly diverse languages and cultures; the majority of its people do not use English as their primary language(s). The utility of English-first AI is limited for the majority of Southeast Asian users, especially in critical sectors like healthcare and education. Through SEACrowd, we have explored the AI landscape in SEA and bridged the gaps in resources, evaluation, and naturalness analysis of AI models in SEA languages. Further, our initiative has nurtured an open-source research community, which will actively continue to add and maintain datasheets and dataloaders, as well as drive AI research and developments in SEA.
@@ Acknowledgments
We would like to thank our amazing contributors: Joshua Spergel, Tiezheng Yu, Parinthapat Pengpun, Bin Wang, Ishan Jindal, Muhammad Satrio, Jipeng Zhang, Bhavish Pahwa, Haryo Akbarianto Wibowo, Hiroki Nomoto, Yohanes Sigit Purnomo W.P., Ahmad Fathan Hidayatullah, Bryan Wilie, Ruhiyah Faradishi Widiaputri, Rafif Rabbani, Fawwaz Mayda, Manoj Khatri, Supryadi Supryadi, Virach Sornlertlamvanich, Pavaris Ruangchutiphophan, Erland Hilman Fuadi, Mega Fransiska, Richardy Sapan, and Camilla Johnine Cosme for their hard work in submitting datasheets and implementing dataloaders for SEACrowd.
This work is supported by the National Research Foundation, Singapore under its AI Singapore Programme; PhD Fellowship Award, the Hong Kong University of Science and Technology; and PF20-43679 Hong Kong PhD Fellowship Scheme, Research Grant Council, Hong Kong. JMI is funded by National University Philippines and the UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI [EP/S023437/1] of the University of Bath. In addition, we would like to express our gratitude to Cohere For AI sfor providing research grants that enabled us to perform our experiments using a commercial baseline, specifically Command-R.
@@
While our work covers nearly 1,000 SEA languages, many dialects, which are considered as belonging to a parent language, are missing from our evaluation benchmark. For instance, for the Malay language, only Standard Malay (zsm) is evaluated, but not other dialects such as Sarawak Malay (zlm-sar). Furthermore, the majority of our datasets also do not contain code-switched texts, which is a common linguistic phenomenon of SEA language usage (Aji et al., 2023). Moreover, the language coverage of different evaluation tasks varies significantly. For instance, text-based tasks cover 34 languages in total, whereas vision-language tasks only cover four languages.
@@
Appendix A Summary of SEACrowd
@@
Addressing the resource gaps and challenges in AI development for Southeast Asian (SEA) languages is essential for our region’s sustainable and prosperous future. The lack of representation of SEA languages in machine learning pre-training models severely impacts their quality. Additionally, the scarcity of high-quality datasets and evaluation tools further hinders progress in AI for SEA languages. The dominance of English-centric training data introduces cultural biases and fails to capture the local values and nuances of SEA cultures. To overcome these obstacles, SEACrowd provides a comprehensive and standardized resource center, along with evaluation tasks, for nearly 1,000 SEA indigenous and non-indigenous languages across various modalities. SEACrowd closes the resource and evaluation gaps, enabling researchers and developers to improve the performance of AI models for SEA languages.
The journey does not end here. Concrete next steps are essential to drive AI advancement in Southeast Asia. Strategic investments in research and development, collaborations with local communities, and efforts toward language preservation are imperative. Governments, industry leaders, and stakeholders must prioritize the development of national and under-resourced local languages to ensure resource equity and promote inclusivity in AI technology. By taking bold actions, such as funding initiatives for data collection and model training, establishing partnerships with local communities, and focusing on language preservation, we can unlock the full potential of Southeast Asian languages. Not only this will spur economic growth but also preserve the region’s rich cultural heritage.
In conclusion, developing AI for Southeast Asian languages is not a mere necessity but an opportunity to create a sustainable and prosperous future. By addressing resource gaps, accurately evaluating models, and fostering inclusive AI development, we can harness the power of SEA languages to drive long-term economic growth while preserving our region’s cultural diversity.
@@ Appendix B Related Work
Open-source and open-science communities play a crucial role in engaging native speakers to curate large-scale multilingual NLP resources. In the past, collaborative efforts have been organized to collect data and train multilingual language models either on a global scale (Workshop et al., 2022; Singh et al., 2024; Üstün et al., 2024) or on a regional level, e.g., Masakhane for African languages (Adelani et al., 2021, 2022b, 2022a, 2023), AI4Bharat for Indian languages (Kakwani et al., 2020; Kumar et al., 2022; Dabre et al., 2022, inter alia), and AmericasNLP for Latin American languages Mager et al. (2021); Ebrahimi et al. (2022). In the SEA region, there have been community-based initiatives, e.g., IndoNLP, PyThaiNLP, and RojakNLP, to study NLP on Indonesian languages (Aji et al., 2022; Wilie et al., 2020; Cahyawijaya et al., 2021, 2023a), Thai language (Phatthiyaphaibun et al., 2023), and the code-switching phenomenon in SEA (Aji et al., 2023; Yong et al., 2023; Winata et al., 2024), respectively.
@@ G.2 Baselines
Table  19, 20, and 21 report the details of baseline models used in SEACrowd evaluation (§3). For each baseline model, we provide information regarding the model size, origin base model, seen languages in the training corpora use, and the URL where the models can be downloaded. In principle, this work does not aim to acquire and fit all available SEA-trained LLMs over the Internet, as this is computationally expensive. Rather, we want to initiate the exploration of select publicly available models to serve as baselines for the evaluation of foundational capabilities on SEA languages through benchmarking on NLU, NLG, speech, and vision tasks aggregated via SEACrowd.
Across the various models explored, as listed in the tables, we prioritized the diversity of model variation in terms of scale, openness, and coverage of SEA languages. In NLP tasks, we covered five language model families for the main experiments, namely English-only, multilingual, regional, and country-specific models. Instruction-tuned LLMs demonstrate the ability to generalize to unseen tasks Wei et al. (2021); Sanh et al. (2021); Ouyang et al. (2022). When these LLMs are based on a multilingual foundation, they have shown proficiency in generalizing across multiple languages Muennighoff et al. (2022); Adilazuarda et al. (2023); Zhang et al. (2023a). For NLU, we compute the weighted F1-score and obtain the answers via log-likelihood for open-source baselines or string matching for commercial baselines.
@@
G.3 Prompts
Tables 22, 23, and 24 describe the handwritten prompt templates used in NLU, NLG, and VL evaluation (§3). For all tasks, we used a zero-shot prompting procedure to serve as the baseline setup. Due to the task complexity and distribution of workload from volunteer contributors with available computing resources, we limited the experiment procedure for some setups to ensure the acquisition of results in line with target release dates. For NLU, we explored three prompt styles for each dataset from core tasks, including commonsense reasoning, question-answering, and NLI. For more challenging tasks requiring more intensive computing power such as NLG and VL, we used only one uniform prompt style, but we also explored prompts translated into SEA languages, i.e., Filipino, Indonesian, Thai, and Vietnamese for VL.
@@ G.3 Prompts
G.4 Evaluation Results
Table 25 and 26 describes the NLU and NLG results per language.
G.5 Language Equity Results
Appendix H Supplementary Details for Translationese Classifier
H.1 Training & Evaluation Data
We manually select and validate the text collection method of each data subset for training and evaluating the translationese classifier, in Tables 27 and 28, respectively. This validation is done by checking the relevant publication, domain, and annotation method. If the texts in the data subsets are a product of machine or human translation, we regard them as translationese. We label data subsets with human-generated texts as natural data.
H.2 Experiments
Classical ML
Encoder LM
We explore fine-tuning encoder-only LM for developing a translationese classifier. We utilize mDeBERTa-v3base model [^25] https://huggingface.co/microsoft/mdeberta-v3-base He et al. (2020, 2022) – a multilingual encoder-only LM – as our backbone. We trained the model with AdamW Loshchilov and Hutter (2019) optimizer using a learning rate of 1e-5, batch size of 256, and warming up steps of 500 for a maximum of 10 epochs. We apply an early stopping of 3 epochs based on the validation accuracy.
@@ Appendix I Supplementary Details for SEA Language Prioritization
Based on the results of the global utility metric Blasi et al. (2022), we provide the top-20 SEA indigenous languages to be prioritized based on their demand (i.e., the number of SEA language speakers) and current utility (Figure 10) or resource availability (Figure 11). [^26] https://github.com/SEACrowd/globalutility While the current utility, also known as the model capability, is relative to the model performance on eng, the resource availability is relative to 500, which is approximately the number of datasets in Korean language available in HuggingFace. The Korean language is chosen as the pivot because it is considered a higher-resource language than most by Joshi et al. (2020).
@@
Table 29-47 present the list of SEA indigenous languages covered by SEACrowd. Information regarding the ISO 639-3 code, language name, region, and population is obtained from Eberhard et al. (2021); Hammarström et al. (2024); Project (2024); Dryer and Haspelmath (2013) and Wikipedia [^27] https://www.wikipedia.org/.
@@
GPT-4 N/A GPT-4 N/A https://openai.com/index/gpt-4/

### v4

@@ 5.2 Conclusion & Future Work
Southeast Asia is home to highly diverse languages and cultures; the majority of its people do not use English as their primary language. The utility of English-first AI is limited for the majority of Southeast Asian users, especially in critical sectors like healthcare and education. Through SEACrowd, we have explored the AI landscape in SEA and bridged the gaps in resources, evaluation, and naturalness analysis of AI models in SEA languages. Further, our initiative has nurtured an open-source research community, which will actively continue to add and maintain datasheets and dataloaders, as well as drive AI research and developments in SEA.
@@ Acknowledgments
We would like to thank our amazing contributors: Joshua Spergel, Tiezheng Yu, Parinthapat Pengpun, Ishan Jindal, Muhammad Satrio, Jipeng Zhang, Bhavish Pahwa, Haryo Akbarianto Wibowo, Hiroki Nomoto, Yohanes Sigit Purnomo W.P., Ahmad Fathan Hidayatullah, Bryan Wilie, Ruhiyah Faradishi Widiaputri, Rafif Rabbani, Fawwaz Mayda, Manoj Khatri, Supryadi Supryadi, Virach Sornlertlamvanich, Pavaris Ruangchutiphophan, Erland Hilman Fuadi, Mega Fransiska, Richardy Sapan, and Camilla Johnine Cosme, for their hard work in submitting datasheets and implementing dataloaders for SEACrowd.
This work is supported by the National Research Foundation, Singapore under its AI Singapore Programme; PhD Fellowship Award, the Hong Kong University of Science and Technology; and PF20-43679 Hong Kong PhD Fellowship Scheme, Research Grant Council, Hong Kong. JMI is funded by National University Philippines and the UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI [EP/S023437/1] of the University of Bath. In addition, we would like to express our gratitude to Cohere For AI for providing research grants that enabled us to perform our experiments using a commercial baseline, specifically Command-R.
@@
While our work covers nearly 1,000 SEA languages, many dialects, which are considered as belonging to a parent language, are missing from our evaluation benchmark. For instance, for the Malay language, only Standard Malay (zsm) is evaluated, but not other dialects such as Sarawak Malay (zlm-sar). Furthermore, the majority of our datasets also do not contain code-switched texts, which is a common linguistic phenomenon of SEA language usage (Aji et al., 2023). Moreover, the language coverage of different evaluation tasks varies significantly. For instance, NLP tasks cover 34 languages in total, whereas VL tasks only cover 4 languages.
@@
Appendix A Key Takeaways of SEACrowd
@@
Key findings include:
Model Performance.
  • LLMs: SEA-specific models, such as AYA-101 and mT0, show strong performance on zero-shot tasks, outperforming English or country-specific models in the region. However, tasks like abstractive QA and summarization reveal limitations in existing models’ ability to handle SEA languages effectively.
  • Speech: Off-the-shelf models like Whisper v3 show competitive ASR performance for major SEA languages but struggle with indigenous languages. In contrast, Seamless M4T v2 offers more balanced results across SEA languages.
  • VLMs: Current VLMs fail to generate high-quality image captions in SEA languages, highlighting the need for more effective multilingual pre-training.
LLM Generation Quality.
SEA language outputs by LLMs are often plagued by translationese, with models like SEA-LION v1 producing natural sentences only 57.71% of the time. Languages like Tagalog, Burmese, and Malay suffer from unnatural generation.
Resource Gaps.
SEACrowd covers 74.9% of SEA languages but reveals a long-tail distribution, where most languages lack comprehensive datasets. SEA languages also face cultural misrepresentation, with 70% of datasets being translations rather than culturally relevant sources.
Prioritizing Development.
Focus should be placed on SEA national languages with significant gaps in naturalness (e.g., Malay, Burmese, Filipino), as well as under-resourced local languages like Javanese and Cebuano.
Collaboration.
Governments, industries, and local communities must invest in R&D, data collection, and open collaboration to address resource equity and improve SEA AI development.
@@ Appendix B Related Work
Open-source and open-science communities play a crucial role in engaging native speakers to curate large-scale multilingual NLP resources. In the past, collaborative efforts have been organized to collect data and train multilingual language models either on a global scale (Workshop et al., 2022; Singh et al., 2024; Üstün et al., 2024) or on a regional level, e.g., Masakhane for African languages (Adelani et al., 2021, 2022b, 2022a, 2023), AI4Bharat for Indian languages (Kakwani et al., 2020; Kumar et al., 2022; Dabre et al., 2022, inter alia), and AmericasNLP for Latin American languages Mager et al. (2021); Ebrahimi et al. (2022).
In the SEA region, there have been community-based initiatives, e.g., IndoNLP, PyThaiNLP, and RojakNLP, to study NLP on Indonesian languages (Aji et al., 2022; Wilie et al., 2020; Cahyawijaya et al., 2021, 2023a), Thai language (Phatthiyaphaibun et al., 2023), and the code-switching phenomenon in SEA (Aji et al., 2023; Yong et al., 2023; Winata et al., 2024), respectively.
@@ G.2 Baselines
Table 20, 21, and 22 report the details of baseline models used in SEACrowd evaluation (§3). For each baseline model, we provide information regarding the model size, origin base model, seen languages in the training corpora use, and the URL where the models can be downloaded. In principle, this work does not aim to acquire and fit all available SEA-trained LLMs over the Internet, as this is computationally expensive. Rather, we want to initiate the exploration of select publicly available models to serve as baselines for the evaluation of foundational capabilities on SEA languages through benchmarking on NLU, NLG, speech, and vision tasks aggregated via SEACrowd.
Across the various models explored, as listed in the tables, we prioritized the diversity of model variation in terms of scale, openness, and coverage of SEA languages. In NLP tasks, we covered 5 LLM groups for the main experiments: English-only, multilingual, regional, and country-specific models. Instruction-tuned LLMs demonstrate the ability to generalize to unseen tasks Wei et al. (2021); Sanh et al. (2021); Ouyang et al. (2022). Some of these LLMs are based on a multilingual foundation, hence their proficiency in generalizing across languages Muennighoff et al. (2022); Adilazuarda et al. (2023); Zhang et al. (2023a). For NLU, we compute the weighted F1-score and obtain the answers via log-likelihood for open-source baselines or string matching for commercial baselines.
@@
G.3 Prompts
Tables 23, 24, and 25 describe the handwritten prompt templates used in NLU, NLG, and VL evaluation (§3). For all tasks, we used a zero-shot prompting procedure to serve as the baseline setup. Due to the task complexity and distribution of workload from volunteer contributors with available computing resources, we limited the experiment procedure for some setups to ensure the acquisition of results in line with target release dates. For NLU, we explored three prompt styles for each dataset from core tasks, including commonsense reasoning, question-answering, and NLI. For more challenging tasks requiring more intensive computing power such as NLG and VL, we used only one uniform prompt style, but we also explored prompts translated into SEA languages, i.e., Filipino, Indonesian, Thai, and Vietnamese for VL.
G.4 Evaluation Results
Table 26 and 27 describes the NLU and NLG results per language.
G.5 Language Equity Results
Appendix H Supplementary Details for Translationese Classifier
H.1 Training & Evaluation Data
We manually select and validate the text collection method of each data subset for training and evaluating the translationese classifier, in Tables 28 and 29, respectively. This validation is done by checking the relevant publication, domain, and annotation method. If the texts in the data subsets are a product of machine or human translation, we regard them as translationese. We label data subsets with human-generated texts as natural data.
H.2 Experiments
Classical ML
Encoder LM
We explore fine-tuning encoder-only LM for developing a translationese classifier. We utilize mDeBERTa-v3base model [^25] https://huggingface.co/microsoft/mdeberta-v3-base He et al. (2020, 2022)–a multilingual encoder-only LM–as our backbone. We train the model with AdamW Loshchilov and Hutter (2019) optimizer using a learning rate of 1e-5, batch size of 256, and warming up steps of 500 for a maximum of 10 epochs. We apply an early stopping of 3 epochs based on the validation accuracy. We show the results in Table 17.
No. Name C. Points
1 Holy Lovenia 549
2 Samuel Cahyawijaya 480
3 Rahmad Mahendra 317
4 Salsabil Maulana Akbar 243
5 Lester James V. Miranda 234
6 Zheng-Xin Yong 164
7 Jennifer Santoso 164
8 Elyanah Aco 158
9 Akhdan Fadhilah 157
10 Jonibek Mansurov 132
11 Fajri Koto 121
12 Joseph Marvin Imperial 118
13 Ruochen Zhang 114
14 Genta Indra Winata 108
15 Onno P. Kampman 107
16 Joel Ruben Antony Moniz 93
17 Muhammad Ravi Shulthan Habibi 92
18 Frederikus Hudi 83
19 Sedrick Keh 81
20 Alham Fikri Aji 80
21 Railey Montalan 78
22 Peerat Limkonchotiwat 72
23 Ryan Ignatius 56
24 Joanito Agili Lopo 50
25 William Nixon 50
26 Börje F. Karlsson 49
27 James Jaya 48
28 Ryandito Diandaru 48
29 Yuze Gao 48
30 William Tjhi 46
31 Patrick Amadeus 46
32 Bin Wang 44
33 Jan Christian Blaise Cruz 43
34 Chenxi Whitehouse 36
35 Ivan Halim Parmonangan 36
36 Maria Khelli 36
37 Sebastian Ruder 35
38 Wenyu Zhang 34
39 Lucky Susanto 33
40 Reynard Adha Ryanda 32
41 Sonny Lazuardi Hermawan 30
42 Dan John Velasco 29
43 Muhammad Dehan Al Kautsar 29
44 Willy Fitra Hendria 29
45 Yasmin Moslem 29
46 Noah Flynn 28
47 Muhammad Farid Adilazuarda 27
48 Haochen Li 27
49 Johanes Lee 27
50 R. Damanhuri 27
51 Shuo Sun 27
52 Muhammad Reza Qorib 26
53 Amirbek Djanibekov 25
54 Wei Qi Leong 25
55 Quyet V. Do 24
56 Niklas Muennighoff 24
57 Tanrada Pansuwan 22
58 Ilham Firdausi Putra 21
59 Yan Xu 21
60 Ayu Purwarianti 20
61 Ngee Chia Tai 20
Table 19: Co-authors ordered by their amount of contribution points.
@@ Appendix I Supplementary Details for SEA Language Prioritization
Based on the results of the global utility metric Blasi et al. (2022), we provide the top-20 SEA indigenous languages to be prioritized based on their demand (i.e., the number of SEA language speakers) and current utility (Figure 10) or resource availability (Figure 11). [^26] https://github.com/SEACrowd/globalutility We use the performance scores of AYA-101 as one of the best-performing models on SEA languages for the current utility. While the current utility, also known as the model capability, is relative to the model performance on eng, the resource availability is relative to 500, which is approximately the number of datasets in Korean language available in HuggingFace. The Korean language is chosen as the pivot because it is considered a higher-resource language than most by Joshi et al. (2020).
@@
Table 30-48 present the list of SEA indigenous languages covered by SEACrowd. Information regarding the ISO 639-3 code, language name, region, and population is obtained from Eberhard et al. (2021); Hammarström et al. (2024); Project (2024); Dryer and Haspelmath (2013) and Wikipedia [^27] https://www.wikipedia.org/.
Appendix L Amount of Contributions by Co-Authors
Table 19 provides a list of co-authors sorted by their amount of contributions in SEACrowd. The full details of their contributions can be seen in our contribution tracking.
@@
GPT-4 N/A GPT-4 N/A https://openai.com/index/gpt-4/. We used turbo-2024-04-09 for NLU and gpt-4o-2024-05-13 for NLG.